OpenStudy (anonymous):

solve with logarithms. 25,000=10000(1.05)^2x

OpenStudy (paxpolaris):

well we can divide both sides by 1000 \[\Large 25=10\left( 1.05 \right)^{2x}\]

OpenStudy (paxpolaris):

\[\large2.5=\left( 1.05 \right)^{2x}\] then we can take the log of both sides...

OpenStudy (paxpolaris):

\[\large \log(2.5)=2x \log(1.05)\] \[\large \implies {\log(2.5)\over \log(1.05)}=2x\\ \large \implies x=\color{green}{\log(2.5)\over 2\cdot \log(1.05)}\]